OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). Question 9 options: 1) -9x + 27 2) -9x + 25 3) -9x + 5 4) -9x - 2

OpenStudy (lalaly):

f(-x+3) = 9(-x+3) -2 = -9x+27-2 = -9x+25

OpenStudy (lalaly):

2)

OpenStudy (anonymous):

This is another good one. Now everywhere you see an x in the f(x) you will write the entire g(x) equation. Then simplify.

OpenStudy (anonymous):

Let f(x) = x - 6 and g(x) = x^2 - x + 4. Find g(f(x)). Question 10 options: 1) x2 - 13x + 46 2) x2 - 11x + 34 3) x2 - 2x - 2 4) x2 - x - 2

OpenStudy (anonymous):

Same thing, replace the variable in g(x) with the expression from f(x). g (x-6) = (x-6)^2 - (x-6) + 4 Now you can expand the binomial squared and simplify.